OpenStudy (laddiusmaximus):

lim x-->8 (x-8/x^2-64)

OpenStudy (rogue):

Can you factor the denominator into anything? Hint: difference of squares

OpenStudy (laddiusmaximus):

?

OpenStudy (rogue):

\[\lim_{x \rightarrow 8} \frac {x-8}{x^2 - 64} = \lim_{x \rightarrow 8} \frac {x-8}{(x-8)(x+8)}\]

OpenStudy (laddiusmaximus):

oh yeah

OpenStudy (rogue):

Your answer should be 1/16. Did you get it?

OpenStudy (laddiusmaximus):

no

OpenStudy (laddiusmaximus):

I know the (x-8) cancels which leaves 1/ x+8

OpenStudy (rogue):

\[\lim_{x \rightarrow 8} \frac {1}{(x+8)} = \frac {1}{16}\]

OpenStudy (rogue):

Yeah, so just plug in 8.

OpenStudy (rogue):

That's what x approaches in the limit.

OpenStudy (laddiusmaximus):

got it! thank you!

OpenStudy (rogue):

Alrighty, no problem :)